OpenStudy (anonymous):

Find the 5th term of the sequence, an=4(3)n - 1. 240 324 972 624

OpenStudy (anonymous):

Just substitute the number 5 into n since the question is asking for 5th term.

OpenStudy (anonymous):

to the both n's

OpenStudy (anonymous):

No, just the n which is in the form of power. (n-1)

OpenStudy (anonymous):

is it 972?

OpenStudy (anonymous):

No, the answer is 324. 4(3)*n-1 is the question right. Substitute 5 into n. 4(3)*5-1=4(3)*4 =4x(3*4) =4x81 =324

myininaya (myininaya):

I don't agree on your order of operations

myininaya (myininaya):

Use the order of opearstions given in the prealgegra books

myininaya (myininaya):

1Exponents 2operations in grouping symbols 3mult/div as it occurs left to right 4add/sub as it occurs left to right

myininaya (myininaya):

Also ask yourself is the equation written correctly because as is I do not see the answer